# Estimating Heterogeneous Causal Effects of High-Dimensional Treatments:

### v4

@@ 1 Introduction
Over the past decade, a number of researchers have exploited modern machine learning algorithms and proposed new methods to estimate heterogeneous treatment effects using experimental data. They include tree-based methods (e.g., Imai and Strauss, 2011; Athey and Imbens, 2016; Wager and Athey, 2018; Hahn, Murray and Carvalho, 2020), regularized regressions (e.g., Imai and Ratkovic, 2013; Tian et al., 2014; Künzel et al., 2019), ensemble methods (e.g., van der Laan and Rose, 2011; Grimmer, Messing and Westwood, 2017), and frameworks that allow for the use of generic machine learning methods (e.g., Chernozhukov et al., 2019; Imai and Li, 2024). This methodological development, however, has largely been confined to settings with a single, binary treatment variable; some exceptions include a time-varying treatment (e.g., Almirall et al., 2014), and a relatively small number of treatments (e.g., Imai and Ratkovic, 2013).
@@
We consider a common setting where researchers wish to use a small number of groups to summarize heterogeneous treatment effects and characterize these groups using several pre-treatment covariates (e.g., Chernozhukov et al., 2019; Imai and Li, 2024). We show that once researchers select the number of groups to be used for summarizing heterogeneous treatment effects, finding the maximally heterogeneous groups in terms of potential outcomes is equivalent to maximizing the likelihood function based on the latent group membership. Furthermore, modeling the conditional probability of an individual’s latent group membership using the moderators of interest yields maximally heterogeneous groups that are predicted well by these moderators.
@@
In the methodological literature on factorial designs and conjoint analysis, researchers have focused on average marginal effects, which represent the average effect of one factor level relative to another level of the same factor averaging over the randomization distribution of the remaining factors (Hainmueller, Hopkins and Yamamoto, 2014; Dasgupta, Pillai and Rubin, 2015). Many empirical researchers use a subgroup analysis to explore how these marginal effects depend on a small number of moderating covariates (e.g., Hainmueller and Hopkins, 2015; Newman and Malhotra, 2019).
@@
The rest of the paper is organized as follows. In Section 2, we discuss the motivating application, which is a conjoint analysis of American citizens’ preferences regarding immigrant features. We also briefly describe a methodological challenge to be addressed. In Section 3, we present our proposed methodology. In Section 4, we show our method performs well in a realistic numerical simulation. In Section 5, we apply this methodology and reanalyze the data from the motivating conjoint analysis. Section 6 concludes with a discussion.
@@ 2.1 The Experimental Design
The attributes of immigrant profiles used in this factorial experiment, with number of levels provided in parentheses, are gender (2), education (7), employment plans (4), job experience (4), profession (11), language skills (4), country of origin (10), reasons for applying (3), and prior trips to the United States (5). For completeness, these factors and their levels are reproduced as Table A1 of Appendix A. In total, there exist over 1.4 million possible profiles, implying more than $2\times 10^{12}$ possible comparisons of two profiles that are possible in the experiment. It is clear that with 1,407 respondents, even though each respondent performs five comparisons, not all possible profiles can be included. Thus, exploring treatment effect heterogeneity requires a methodological development that goes beyond the models used previously in the causal inference literature for binary treatments.
@@ 3.2 General Framework
$\delta_{j}(l,l^{\prime})\ =\ \mathbb{E}[Y_{i}(T_{ij}=l,\bm{T}_{i,-j})-Y_{i}(T_{ij}=l^{\prime},\bm{T}_{i,-j})],$ (1)
where $l\neq l^{\prime}\in\mathcal{T}_{j}$ with $\mathcal{T}_{j}$ representing the support of the randomization distribution for $T_{j}$. The expectation in Equation (1) is taken over the distribution of the other factors $\bm{T}_{i,-j}$ as well as the random sampling of units from the population. Thus, the AMCE averages over two sources of causal heterogeneity — heterogeneity across treatment combinations and across units. Different treatment combinations may have distinct impacts on units with varying characteristics. Our goal is to model these potentially complex heterogeneous treatment effects using an interpretable model.
We propose to model heterogeneous treatment effects based on $K$ distinct treatment effect patterns where $K\geq 2$ is chosen by a researcher, based on their desired granularity of heterogeneity. This approach, which is based on a fixed number of subgroups to characterize treatment effect heterogeneity, is commonly used by empirical researchers. Others have studied various methodological aspects of this approach albeit in the context of binary treatment (Chernozhukov et al., 2019; Imai and Li, 2024).
@@
Proposition 1 (Finding maximally heterogeneous groups)
@@
$\operatornamewithlimits{argmax}_{\mathcal{Z}}\left\{\sum_{k=1}^{K}\sum_{i=1}^{N}\mathbf{1}\{Z_{i}=k\}\mathrm{KL}\left(\hat{\zeta}_{k}(\bm{T}_{i};\mathcal{Z})\|\widehat{\overline{Y}}(\bm{T}_{i})\right)\right\}$ (2)
@@
where $Y_{i}$ is binary, the KL divergence of two Bernoulli distributions with means $\mu_{1}$ and $\mu_{2}$ is given by $\mathrm{KL}\left(\mu_{1}\|\mu_{2}\right)\ =\ \mu_{1}\log\mu_{1}/\mu_{2}+(1-\mu_{1})\log(1-\mu_{1})/(1-\mu_{2})$, and $\{\hat{\zeta}_{k}(\bm{t};\mathcal{Z})\}_{k=1}^{K}$ denotes the maximizers of the right hand side of Equation (2) given $\mathcal{Z}$.
Appendix C provides a proof of a more general result for the natural exponential family distributions (see Chi, Chi and Baraniuk, 2016, for a similar result in the Gaussian case). The log-likelihood formulation is equivalent to the classification maximum likelihood approach in mixture modeling (McLachlan, 1982).
@@
Proposition 2 (Finding maximally heterogeneous groups with moderators)
@@
$\operatornamewithlimits{argmax}_{\mathcal{Z}}\left\{\sum_{k=1}^{K}\sum_{i=1}^{N}\mathbf{1}\{Z_{i}=k\}\mathrm{KL}\left(\hat{\zeta}_{k}(\bm{T}_{i};\mathcal{Z})\|\widehat{\overline{Y}}(\bm{T}_{i})\right)-\sum_{i=1}^{N}H(\{\hat{\pi}_{k}(\bm{X}_{i};\mathcal{Z})\}_{k=1}^{K})\right\}$ (3)
@@
where $H(\{p_{k}\}_{k=1}^{K})=-\sum_{k=1}^{K}p_{k}\log p_{k}$ (by convention, if $p_{k}=0$, then $p_{k}\log p_{k}=0$) is the entropy, and $\hat{\pi}_{k}(\bm{x};\mathcal{Z})=\sum_{i=1}^{N}\bm{1}\{Z_{i}=k,\bm{X}_{i}=\bm{x}\}/\sum_{i=1}^{N}\bm{1}\{\bm{X}_{i}=\bm{x}\}$ and $\hat{\zeta}_{k}(\bm{t};\mathcal{Z})$ are the maximizers of the log-likelihood function of the right hand side of Equation (3) given $\mathcal{Z}$.
@@
Proof is given in Appendix D. Since the entropy $H(\{\hat{\pi}_{k}(\bm{x};\mathcal{Z})\}_{k=1}^{K})$ is maximized when $\hat{\pi}_{k}(\bm{x})=1/K$, Proposition 2 shows that adding a group membership model based on moderators encourages finding groups whose memberships are well predicted by the moderators.
Directly optimizing Equation (2) or (3) over $\mathcal{Z}$ is difficult given its high-dimensionality. Indeed, the resulting estimator is likely to be inconsistent due to the incidental parameter problem (Bryant and Williamson, 1978). To address this problem, we take a Bayesian perspective by interpreting the objective function of the right hand side of Equation (3) as the log-posterior that consists of the log-likelihood and the log-prior over the latent group memberships $\mathcal{Z}$ (McLachlan, 1982). By marginalizing (instead of optimizing) over $\mathcal{Z}$ and adding priors on model parameters $\{\zeta_{k},\pi_{k}\}_{k=1}^{K}$, the optimization problem becomes equivalent to finding the posterior mode of $p(\{\zeta_{k},\pi_{k}\}_{k=1}^{K}\mid\{Y_{i},\bm{T}_{i},\bm{X}_{i}\}_{i=1}^{N})$.
@@
The above results justify the use of a finite mixture model known as a mixture of experts for heterogeneous effect estimation (e.g., Gormley and Frühwirth-Schnatter, 2019; Gupta and Chintagunta, 1994). We stress that the motivation for the Bayesian perspective is mainly computational. Importantly, our framework does not assume a specific data generating process. Instead, we have shown that given the number of groups and appropriate prior distributions, researchers can find maximally heterogeneous groups by fitting a mixture of experts.
@@ 3.3 Model Specification
Since $\mathcal{T}$ is high-dimensional, many treatment combinations are unobserved with a typical sample size. Thus, nonparametric estimation is not applicable. We, therefore, model $\zeta_{k}(\bm{t})$ using a regularized logistic regression where ANOVA-style sum-to-zero constraint is imposed separately for each factor to facilitate merging of different levels within each factor. This modeling strategy identifies a relatively small number of treatment combinations while avoiding the specification of a baseline level for each factor (Egami and Imai, 2019). The interpretation of $\zeta_{k}(\bm{t})$ under this model is still the average of potential outcome under treatment $\bm{t}$ in group $k$. Note that we do not assume homogeneity of outcomes or effects within each group.
@@
$\zeta_{k}(\bm{T}_{i})\ =\ \frac{\exp(\psi_{k}(\bm{T}_{i}))}{1+\exp(\psi_{k}(\bm{T}_{i}))},\quad\text{and}\quad\pi_{k}(\bm{X}_{i})=\frac{\exp(\bm{X}_{i}^{\top}\bm{\phi}_{k})}{\sum_{k^{\prime}=1}^{K}\exp(\bm{X}_{i}^{\top}\bm{\phi}_{k^{\prime}})},$ (4)
@@
for each $k=1,2,\ldots,K$ where $\tilde{\bm{T}}_{i}$ is the vector of indicators, $\mathbf{1}\{T_{ij}=l\}$ and $\mathbf{1}\{T_{ij}=l,T_{ij^{\prime}}=l^{\prime}\}$, and $\bm{\beta}_{k}$ is a stacked column vector containing all coefficients for group $k$. Inclusion of higher-order interactions is straightforward and hence is omitted for notational simplicity.
@@
$\sum_{l=0}^{L_{j}-1}\beta^{j}_{kl}=0,\quad\text{and}\quad\sum_{l=0}^{L_{j}-1}\beta^{jj^{\prime}}_{kll^{\prime}}=\sum_{l^{\prime}=0}^{L_{j^{\prime}}-1}\beta^{jj^{\prime}}_{kll^{\prime}}=0,$ (5)
@@
where each row of $\bm{C}^{\top}\bm{\beta}_{k}$ corresponds to one of the constraints given in Equation (5).
3.4 Sparsity-induced Prior

### v5

@@ 1 Introduction
Over the past decade, a number of researchers have exploited modern machine learning algorithms and proposed new methods to estimate heterogeneous treatment effects using experimental data. They include tree-based methods (e.g., Imai and Strauss, 2011; Athey and Imbens, 2016; Wager and Athey, 2018; Hahn, Murray and Carvalho, 2020), regularized regressions (e.g., Imai and Ratkovic, 2013; Tian et al., 2014; Künzel et al., 2019), ensemble methods (e.g., van der Laan and Rose, 2011; Grimmer, Messing and Westwood, 2017), and frameworks that allow for the use of generic machine learning methods (e.g., Chernozhukov et al., 2019; Imai and Li, 2025). This methodological development, however, has largely been confined to settings with a single, binary treatment variable; some exceptions include a time-varying treatment (e.g., Almirall et al., 2014), and a relatively small number of treatments (e.g., Imai and Ratkovic, 2013).
@@
We consider a common setting where researchers wish to use a small number of groups to summarize heterogeneous treatment effects and characterize these groups using several pre-treatment covariates (e.g., Chernozhukov et al., 2019; Imai and Li, 2025). We show that once researchers select the number of groups to be used for summarizing heterogeneous treatment effects, finding the maximally heterogeneous groups in terms of potential outcomes is equivalent to maximizing the likelihood function based on the latent group membership. Furthermore, modeling the conditional probability of an individual’s latent group membership using the moderators of interest yields maximally heterogeneous groups that are predicted well by these moderators.
@@
In the methodological literature on factorial designs and conjoint analysis, researchers have focused on average marginal effects, which represent the average effect of one factor level relative to another level of the same factor averaging over the randomization distribution of the remaining factors (Hainmueller, Hopkins and Yamamoto, 2014; Dasgupta, Pillai and Rubin, 2015). Many empirical researchers use subgroup analysis to explore how these marginal effects depend on a small number of moderating covariates (e.g., Hainmueller and Hopkins, 2015; Newman and Malhotra, 2019).
@@
The rest of the paper is organized as follows. In Section 2, we discuss the motivating application, which is a conjoint analysis of American citizens’ preferences regarding immigrant features. We also briefly describe a methodological challenge to be addressed. In Section 3, we present our proposed methodology. In Section 4, we show our method performs well in a realistic numerical simulation. In Section 5, we apply this methodology and reanalyze the data from the motivating conjoint analysis. Section 6 concludes with a discussion. The R package FactorHet (Goplerud, Pashley and Imai, 2025) can be used to implement our methodology and Goplerud, Imai and Pashley (2025) provides replication code for our application and simulations.
@@ 2.1 The Experimental Design
The attributes of immigrant profiles used in this factorial experiment, with number of levels provided in parentheses, are gender (2), education (7), employment plans (4), job experience (4), profession (11), language skills (4), country of origin (10), reasons for applying (3), and prior trips to the United States (5). For completeness, these factors and their levels are reproduced as Table A1 of the Supplementary Material (Goplerud, Imai and Pashley, in-press). In total, there exist over 1.4 million possible profiles, implying more than $2\times 10^{12}$ possible comparisons of two profiles that are possible in the experiment. It is clear that with 1,407 respondents, even though each respondent performs five comparisons, not all possible profiles can be included. Thus, exploring treatment effect heterogeneity requires a methodological development that goes beyond the models used previously in the causal inference literature for binary treatments.
@@ 3.2 General Framework
$\delta_{j}(l,l^{\prime})\ =\ \mathbb{E}[Y_{i}(T_{ij}=l,\bm{T}_{i,-j})-Y_{i}(T_{ij}=l^{\prime},\bm{T}_{i,-j})],$ (3.1)
where $l\neq l^{\prime}\in\mathcal{T}_{j}$ with $\mathcal{T}_{j}$ representing the support of the randomization distribution for $T_{j}$. The expectation in Equation (3.1) is taken over the distribution of the other factors $\bm{T}_{i,-j}$ as well as the random sampling of units from the population. Thus, the AMCE averages over two sources of causal heterogeneity—heterogeneity across treatment combinations and across units. Different treatment combinations may have distinct impacts on units with varying characteristics. Our goal is to model these potentially complex heterogeneous treatment effects using an interpretable model.
We propose to model heterogeneous treatment effects based on $K$ distinct treatment effect patterns where $K\geq 2$ is chosen by a researcher, based on their desired granularity of heterogeneity. This approach, which is based on a fixed number of subgroups to characterize treatment effect heterogeneity, is commonly used by empirical researchers. Others have studied various methodological aspects of this approach albeit in the context of binary treatment (Chernozhukov et al., 2019; Imai and Li, 2025).
@@
Proposition 1 (Finding maximally heterogeneous groups).
@@
$\operatornamewithlimits{argmax}_{\mathcal{Z}}\left\{\sum_{k=1}^{K}\sum_{i=1}^{N}\mathbf{1}\{Z_{i}=k\}\mathrm{KL}\left(\hat{\zeta}_{k}(\bm{T}_{i};\mathcal{Z})\|\widehat{\overline{Y}}(\bm{T}_{i})\right)\right\}$ (3.2)
@@
where $Y_{i}$ is binary, the KL divergence of two Bernoulli distributions with means $\mu_{1}$ and $\mu_{2}$ is given by $\mathrm{KL}\left(\mu_{1}\|\mu_{2}\right)\ =\ \mu_{1}\log\mu_{1}/\mu_{2}+(1-\mu_{1})\log(1-\mu_{1})/(1-\mu_{2})$, and $\{\hat{\zeta}_{k}(\bm{t};\mathcal{Z})\}_{k=1}^{K}$ denotes the maximizers of the right hand side of Equation (3.2) given $\mathcal{Z}$.
Section C of the Supplementary Material provides a proof of a more general result for the natural exponential family distributions (see Chi, Chi and Baraniuk, 2016, for a similar result in the Gaussian case). The log-likelihood formulation is equivalent to the classification maximum likelihood approach in mixture modeling (McLachlan, 1982).
@@
Proposition 2 (Finding maximally heterogeneous groups with moderators).
@@
$\operatornamewithlimits{argmax}_{\mathcal{Z}}\left\{\sum_{k=1}^{K}\sum_{i=1}^{N}\mathbf{1}\{Z_{i}=k\}\mathrm{KL}\left(\hat{\zeta}_{k}(\bm{T}_{i};\mathcal{Z})\|\widehat{\overline{Y}}(\bm{T}_{i})\right)-\sum_{i=1}^{N}H(\{\hat{\pi}_{k}(\bm{X}_{i};\mathcal{Z})\}_{k=1}^{K})\right\}$ (3.3)
@@
where $H(\{p_{k}\}_{k=1}^{K})=-\sum_{k=1}^{K}p_{k}\log p_{k}$ (by convention, if $p_{k}=0$, then $p_{k}\log p_{k}=0$) is the entropy, and $\hat{\pi}_{k}(\bm{x};\mathcal{Z})=\sum_{i=1}^{N}\bm{1}\{Z_{i}=k,\bm{X}_{i}=\bm{x}\}/\sum_{i=1}^{N}\bm{1}\{\bm{X}_{i}=\bm{x}\}$ and $\hat{\zeta}_{k}(\bm{t};\mathcal{Z})$ are the maximizers of the log-likelihood function of the right hand side of Equation (3.3) given $\mathcal{Z}$.
@@
Proof is given in Section D of the Supplementary Material. Since the entropy $H(\{\hat{\pi}_{k}(\bm{x};\mathcal{Z})\}_{k=1}^{K})$ is maximized when $\hat{\pi}_{k}(\bm{x})=1/K$, Proposition 2 shows that adding a group membership model based on moderators encourages finding groups whose memberships are well predicted by the moderators.
Direct optimization of Equations (3.2) and (3.3) over $\mathcal{Z}$ has been studied under the name of “classification maximum likelihood” in the literature on mixture models (McLachlan, 1982). For completeness, Section G.3 of the Supplementary Material provides an estimation algorithm for this approach, which modifies the proposed algorithm described in Section 3.5. Unfortunately, the classification maximum likelihood approach suffers from the incidental parameter problem because the cardinality of $\mathcal{Z}$ increases with the sample size $N$, leading to an asymptotic bias and inconsistency (Bryant and Williamson, 1978).
To address this problem, a dominant approach in the literature is Bayesian, treating the right hand side of Equation (3.3) as a log-posterior that consists of a log-likelihood and a log-prior over $\mathcal{Z}$, i.e., $\mathrm{Pr}(Z_{i}=k\mid\bm{X}_{i})=\pi_{k}(\bm{X}_{i})$. By marginalizing out $\mathcal{Z}$, we avoid the incidental parameter problem, yielding the objective function known as a mixture maximum likelihood (McLachlan, 1982).
The model is called “mixture-of-experts” when $\pi_{k}$ depends on $\bm{X}_{i}$ (Gormley and Frühwirth-Schnatter, 2019) with the following objective function,
$\{\hat{\zeta}_{k},\hat{\pi}_{k}\}_{k=1}^{K}=\operatornamewithlimits{argmax}_{\{\zeta_{k},\pi_{k}\}_{k=1}^{K}}\sum_{i=1}^{N}\log\left[\sum_{k=1}^{K}\pi_{k}(\bm{X}_{i})\zeta_{k}(\bm{T}_{i})^{Y_{i}}\{1-\zeta_{k}(\bm{T}_{i})\}^{1-Y_{i}}\right].$ (3.4)
While this setup no longer appears to provide a direct characterization of the optimal groups, Proposition 3 shows that a mixture-of-experts model finds maximally heterogeneous groups as in Proposition 2 but with an additional penalty that encourages less extreme posterior probabilities of group memberships.
Proposition 3 (Finding maximally heterogeneous groups with a mixture of experts).
Maximizing the likelihood function under a mixture-of-experts model is equivalent to finding maximally heterogeneous groups as in Proposition 2 with an additional penalty. That is, the following equality holds for any $\mathcal{Z}$,
$\operatornamewithlimits{argmax}_{\zeta,\pi}\sum_{i=1}^{N}\log\left[\sum_{k=1}^{K}\pi_{k}(\bm{X}_{i})\zeta_{k}(\bm{T}_{i})^{Y_{i}}\{1-\zeta_{k}(\bm{T}_{i})\}^{1-Y_{i}}\right]$
$=$ $\operatornamewithlimits{argmax}_{\zeta,\pi}\sum_{i=1}^{N}\sum_{k=1}^{K}\mathbf{1}\{Z_{i}=k\}\left[Y_{i}\log\zeta_{k}(\bm{T}_{i})+(1-Y_{i})\log\{1-\zeta_{k}(\bm{T}_{i})\}\right.$
$\left.\hskip 72.26999pt+\log\pi_{k}(\bm{X}_{i})-\log\tilde{\pi}_{k}(\bm{X}_{i},Y_{i},\bm{T}_{i};\{\zeta_{k^{\prime}},\pi_{k^{\prime}}\}_{k^{\prime}=1}^{K})\right],$
where
$\tilde{\pi}_{k}(\bm{X}_{i},Y_{i},\bm{T}_{i};\{\zeta_{k^{\prime}},\pi_{k^{\prime}}\}_{k^{\prime}=1}^{K})\ =$ $\mathrm{Pr}(Z_{i}=k\mid Y_{i},\bm{T}_{i},\bm{X}_{i},\{\zeta_{k^{\prime}},\pi_{k^{\prime}}\}_{k^{\prime}=1}^{K})$
[…]
is the posterior membership probability for group $k$.
Proof of the proposition directly follows from a well-known identity (e.g., Celeux, Frühwirth-Schnatter and Robert, 2019), and hence is omitted. The equality in Proposition 3 holds for any group membership $\mathcal{Z}$, including its maximum-a-posteriori (MAP) estimate, i.e.,
$\hat{Z}_{i}=\operatornamewithlimits{argmax}_{k}\tilde{\pi}_{k}(\bm{X}_{i},Y_{i},\bm{T}_{i};\{\hat{\zeta}_{k^{\prime}},\hat{\pi}_{k^{\prime}}\}_{k^{\prime}=1}^{K})$. Thus, our proposed model can be seen as finding maximally heterogeneous groups while imposing a penalty that encourages finding groups that are well predicted by the moderators $\bm{X}_{i}$ but with less extreme group membership probabilities based on the data.
@@
All together, our results provide a justification for using a mixture-of-experts model for heterogeneous effect estimation under the settings with high-dimensional treatments. We emphasize that a primary motivation for the use of Bayesian approach is to resolve the incidental parameter problem with classification maximum likelihood. Importantly, the results above do not assume a specific data generating process. Instead, we have shown that given the number of groups and appropriate prior distributions, researchers can find maximally heterogeneous groups by fitting a mixture-of-experts model.
@@ 3.3 Model Specification
Since $\mathcal{T}$ is high-dimensional, many treatment combinations are unobserved with a typical sample size. Thus, nonparametric estimation is not applicable. We, therefore, model $\zeta_{k}(\bm{t})$ using a regularized logistic regression where an ANOVA-style sum-to-zero constraint is imposed separately for each factor to facilitate merging of different levels within each factor. This modeling strategy identifies a relatively small number of treatment combinations while avoiding the specification of a baseline level for each factor (Egami and Imai, 2019). The interpretation of $\zeta_{k}(\bm{t})$ under this model is still the average of potential outcome under treatment $\bm{t}$ in group $k$. Note that we do not assume homogeneity of outcomes or effects within each group.
@@
$\zeta_{k}(\bm{T}_{i})\ =\ \frac{\exp(\psi_{k}(\bm{T}_{i}))}{1+\exp(\psi_{k}(\bm{T}_{i}))},\quad\text{and}\quad\pi_{k}(\bm{X}_{i})=\frac{\exp(\bm{X}_{i}^{\top}\bm{\phi}_{k})}{\sum_{k^{\prime}=1}^{K}\exp(\bm{X}_{i}^{\top}\bm{\phi}_{k^{\prime}})},$ (3.5)
@@
for each $k=1,2,\ldots,K$ where $\tilde{\bm{T}}_{i}$ is the vector of indicators, $\mathbf{1}\{T_{ij}=l\}$ and $\mathbf{1}\{T_{ij}=l,T_{ij^{\prime}}=l^{\prime}\}$, and $\bm{\beta}_{k}$ is a stacked column vector containing all coefficients for group $k$. Inclusion of higher-order interactions is straightforward (see Section E of the Supplementary Material) and hence is omitted in the main paper for notational simplicity.
@@
$\sum_{l=0}^{L_{j}-1}\beta^{j}_{kl}=0,\quad\text{and}\quad\sum_{l=0}^{L_{j}-1}\beta^{jj^{\prime}}_{kll^{\prime}}=\sum_{l^{\prime}=0}^{L_{j^{\prime}}-1}\beta^{jj^{\prime}}_{kll^{\prime}}=0,$ (3.6)
@@
where each row of $\bm{C}^{\top}\bm{\beta}_{k}$ corresponds to one of the constraints given in Equation (3.6).
3.4 Sparsity-inducing Prior